Mio is 18 years older than Yui. Fourteen years ago the sum of their ages was 20. How old are they now?
Answer by richwmiller(17219)

yui=(20-(-28)-18)/2
yui=(30)/2
yui=15
mio=15+18
mio=33
yui=15 mio= 33